I can't figure out how to solve this system of equations with the substitution method. I have tried combining the terms and distributing but I keep getting confused with the process. can you help me?
x=y+6
x+2y=6

Answer by Alan3354(69443)

x=y+6
x+2y=6
------
Sub for x into the 2nd eqn
(y+6) + 2y = 6
3y+6 = 6
3y = 0
y = 0
------
x = 6